### v1

# Annotator-Centric Active Learning for Subjective NLP Tasks

Abstract
To accurately capture the variability in human judgments for subjective NLP tasks, incorporating a wide range of perspectives in the annotation process is crucial. Active Learning (AL) addresses the high costs of collecting human annotations by strategically annotating the most informative samples. We introduce Annotator-Centric Active Learning (ACAL), which incorporates an annotator selection strategy following data sampling. Our objective is two-fold: (1) to efficiently approximate the full diversity of human judgments, and (2) to assess model performance using annotator-centric metrics, which emphasize minority perspectives over a majority. We experiment with multiple annotator selection strategies across seven subjective NLP tasks, employing both traditional and novel, human-centered evaluation metrics. Our findings indicate that ACAL improves data efficiency and excels in annotator-centric performance evaluations. However, its success depends on the availability of a sufficiently large and diverse pool of annotators to sample from.
1 Introduction
A challenging aspect of natural language understanding (NLU) is the variability of human judgment and interpretation in subjective tasks (e.g., hate speech detection) (Plank, 2022). While humans can navigate subjectivity naturally, most machine learning methods are insensitive to individual differences (Sandri et al., 2023) and underrepresented perspectives (van der Meer et al., 2024).
Modern NLU approaches are commonly trained and tested on annotated datasets. In a subjective task, each data sample is typically labeled by a set of annotators, and differences in annotation are reconciled through aggregation techniques (e.g., majority voting), resulting in a single “gold label” (Uma et al., 2021). This approach, though effective for training ML algorithms, neglects the labels of minorities, which becomes problematic, especially, in the case of sensitive subjective tasks.
Subjectivity has been addressed by modeling the full distribution of annotations for each data sample as opposed to aggregating them (Plank, 2022). However, resources for such approaches are scarce, as most datasets do not (yet) make fine-grained annotation details available (Cabitza et al., 2023), and representing a full range of perspectives is contingent on obtaining annotations from a diverse crowd (Bakker et al., 2022).
One way of accounting for a limited annotation budget is to use Active Learning (Settles, 2012, AL). Given a pool of unannotated data samples, AL employs a sample selection strategy to select maximally informative samples for training, retrieving the corresponding annotations from a ground truth oracle (e.g., a single human expert). However, in subjective tasks there is no such oracle, instead we rely on a set of available annotators. Given this practical constraint, we argue that informativeness for AL manifests in both samples and annotations, as the model should also be guided to reflect the distribution of annotations. Demanding all available annotators to annotate all selected samples would provide a truthful representation of the annotation distribution, but is often unfeasible, especially if the pool of annotators is large. Thus, deciding which annotator(s) should annotate the selected samples is as critical as selecting which samples to annotate.
We introduce Annotator-Centric Active Learning (ACAL) to account for annotation diversity in subjective tasks. In ACAL, the sample selection strategy of traditional AL is followed by an annotator selection strategy as Figure 1 shows. For each data sample selected through the sample selection strategy, the annotator selection strategy selects an annotator from the available annotators. We make the following contributions: (1) We create ACAL, extending the AL approach to optimize for diversity among annotators when learning soft labels in subjective tasks. (2) We introduce a suite of annotator-centric evaluation metrics to measure both representativeness and diversity. (3) We demonstrate our approach’s effectiveness on three diverse datasets with subjective tasks—hate speech detection, moral value classification, and safety judgments.
Our experiments show that ACAL works to better approximate the distribution of human judgments with a lower annotation budget. However, this effectiveness requires a large pool of diverse annotators, as is the case for one of our datasets. In other cases, the differences between ACAL and traditional AL become smaller. Through our annotator-centric evaluation, we show that task agreement and the number of available annotations both influence the effectiveness of ACAL, hinting at a direct trade-off between learning from a majority versus being sensitive to minority annotations.
2 Related work
We review related works on annotator disagreement and active learning. Our work is novel in combining these fields to (1) represent annotation distributions through soft labels, (2) incorporate annotator selection strategies in the active learning loop, and (3) evaluate with annotator-centric metrics next to traditional evaluation.
2.1 Learning with annotator disagreement
Modeling annotator disagreement is garnering increasing attention (Aroyo and Welty, 2015; Uma et al., 2021; Plank, 2022; Cabitza et al., 2023). For instance, some aggregation methods can lead to a fairer representation than simple majority (Hovy et al., 2013; Tao et al., 2018). Alternatively, the full annotation distribution can be modeled using soft labels (e.g., Peterson et al., 2019; Müller et al., 2019; Fornaciari et al., 2021; Collins et al., 2022). Other approaches leverage annotator-specific information, e.g., by including individual classification heads per annotator (Davani et al., 2022), embedding annotator-specific behavior (Mokhberian et al., 2023), or encoding the annotator’s socio-demographic information (Beck et al., 2023).
Yet, representing annotator diversity remains challenging. Standard calibration metrics under human label variation may be unsuitable, especially when the variation is high (Baan et al., 2022). Trade-offs ought to be made between collecting more samples or more annotations Gruber et al. (2024). Further, solely measuring differences among sociodemographic traits is not sufficient to fully capture opinion diversity (Orlikowski et al., 2023). To this end, we represent diversity based on which annotators have annotated, what they annotated, and how they have annotated. We experiment with different annotator selection strategies to reveal what aspects impact task performance and annotation budget.
2.2 Active Learning
AL enables a supervised learning model to achieve high performance with a few training examples if chosen judiciously Settles (2012). In a typical AL scenario, a vast collection of unlabeled data is available, and an oracle (e.g., a human expert) can be asked to annotate this unlabeled data. A sampling strategy is employed to iteratively (and smartly) select the next batch of unlabeled data for annotation by the oracle Ren et al. (2021).
AL has found widespread application in the field of NLP Zhang et al. (2022). Two main strategies are employed, either by selecting the unlabeled samples on which the model prediction is most uncertain (Zhang et al., 2017), or by selecting samples that are most representative of the unlabeled dataset (Erdmann et al., 2019; Zhao et al., 2020).
The combination of AL and annotator diversity is a novel direction that has not garnered much attention yet. Existing work proposes to align model uncertainty with annotator uncertainty (Baumler et al., 2023), whereas others adapt annotator-specific classification heads in AL settings (Wang and Plank, 2023), or select texts to annotate based on annotator preferences (Kanclerz et al., 2023).
Existing methods ignore a crucial part of learning with human variation: the diversity among annotators. We focus on which annotators should annotate, such that it best informs us about the underlying label diversity.
3 Method
First, we define the soft-label prediction task we use to train a supervised model. Then, we introduce the traditional AL and the novel ACAL approaches.
3.1 Soft-Label prediction
Consider a dataset composed of triples , where is a data sample (i.e., a piece of text) and is the class label assigned by annotator . The multiple labels assigned to a sample by the different annotators are usually combined into an aggregated label . For training with soft labels, the aggregation typically takes the form of maximum likelihood estimation Uma et al. (2021):
In our experiments, We use a passive learning approach that uses all available to train a model with cross-entropy loss as a baseline.
3.2 Active Learning
AL imposes a sampling technique for inputs , such that the most informative sample(s) are picked for learning. In a typical AL approach, a set of unlabelled data points is available. At every iteration, a sample selection strategy selects samples to be annotated by an oracle that provides the ground truth label distribution . The selected samples and annotations are added to the labeled data , with which the model is trained. Alg. 1 provides an overview of the procedure. In our sample selection strategies, a batch of data of a given size is queried at each iteration. In our experiments, we compare the following strategies:
- Random ()
selects a samples uniformly at random from .
- Uncertainty ()
predicts a distribution over class labels with for each . Select the samples with the highest prediction entropy (i.e., the samples on which the model is most uncertain).
3.3 Annotator-Centric Active Learning
The ACAL approach builds on the AL approach. In contrast to AL, which retrieves an aggregated annotation , ACAL employs an annotator selection strategy to select one annotator and their annotation for each selected data point . Alg. 2 describes the ACAL approach.
Dataset | Task (dimension) | Num. Samples | Num. Annotators | Num. Annotations | Avg. Annotations per item |
DICES | Safety Judgment | 990 | 172 | 72,103 | 72.83 |
MFTC | Morality (care) | 8434 | 23 | 31310 | 3.71 |
MFTC | Morality (loyalty) | 3288 | 23 | 12803 | 3.89 |
MFTC | Morality (betrayal) | 12546 | 23 | 47002 | 3.75 |
MHS | Hate Speech (dehumanize, genocide, respect) | 17282 | 7807 | 57980 | 3.35 |
We propose annotator selection strategies that include annotations from diverse annotators. The strategies vary in the type of information used to represent differences between annotators, and include what or how the annotators have annotated thus far. We test the following strategies:
- Random ()
selects one random annotator .
- Label Minority ()
considers only the labels that annotators have assigned. Given a new sample , selects the available annotator that has the largest bias toward the minority label compared to the other available annotators, i.e., who has annotated other samples with the minority label the most. The minority label is selected as the class with the smallest annotation count in the available dataset thus far.
- Semantic Diversity ()
considers only information on what each annotator has annotated so far (i.e., the samples that they have annotated). Given a new sample selected through , selects the available annotator for whom is semantically the most different from what the annotator has labeled so far. To measure this difference for an annotator , we employ a sentence embedding model to measure the cosine distance between the embeddings of and embeddings of all the samples annotated by . We then take the average of all semantic similarities. The annotator with the lowest average similarity score is selected.
- Representation Diversity ()
selects the annotator that has the lowest similarity with the other annotators available for that item. We create a simple representation for each annotator based on the items together with the respective label that they have annotated, followed by computing the pair-wise cosine similarity between all annotators.
4 Experimental Setup
We describe the experimental setup for the comparisons between ACAL strategies. In all our experiments, we employ a TinyBERT model (Jiao et al., 2019) to reduce the number of trainable parameters. Appendix A includes a detailed overview of the computational setup and hyperparameters. We will provide our codebase upon publication.
4.1 Datasets
Table 1 introduces the three datasets that we use, with variation in domain, annotation task (in italics), annotator count, and annotations per instance.
The DICES Corpus (Aroyo et al., 2023) is composed of 990 conversations with an LLM where 172 annotators provided judgments on whether a generated response can be deemed safe (3-way judgments: yes, no, unsure). We perform a multi-class classification with the scores.
The MFTC Corpus Hoover et al. (2020) is composed of 35K tweets that 23 annotators annotated with any of the 10 moral elements from the Moral Foundation Theory Graham et al. (2013). We select the elements of loyalty (lowest annotation count), care (medium count), and betrayal (highest count) and perform three binary classifications to predict the presence of the respective elements. As most tweets were labeled non-moral (i.e., with no moral element), we balanced the datasets by subsampling the non-moral class.
The MHS Corpus (Sachdeva et al., 2022) consists of 50K social media comments on which 8K annotators judged three hate speech aspects—dehumanize (low inter-rater agreement), respect (medium agreement), and genocide (high agreement)—on a 5-point Likert scale. We perform a multi-class classification with the annotated Likert scores for each task.
The datasets and tasks differ in the entropy scores over annotations (Appendix A.5). DICES and MHS generally have medium normalized entropy scores (most lie between , whereas the MFTC entropy scores are highly polarized.
4.2 Training procedure
We test the annotator selection strategies proposed in Section 3.3 by comparing all possible combinations of the two different sample selection strategies ( and ) with the annotator selection strategies (, , , and ). At each round, we use to select unique samples from the unlabeled data pool . We empirically select to be the smallest between 5% of the number of available annotations and the number of unique samples in the training set. For each selected sample , we use to select one annotator and retrieve their annotation .
To populate the annotation history for the annotation selection strategies, we perform a single warmup round with randomly selected annotations before starting the ACAL iterations (Zhang et al., 2022). We report our training progress results on a validation set with 3-fold cross-validation, showing the average to account for stability across random data splits (into 80% train, 10% validation, and 10% test) and initialization. Then, we select the model iteration that led to the best performance (according to ) on the validation set and evaluate it using a separate test set.
We compare our work with traditional Oracle-based AL approaches ( and ), which use the data sampling strategies but obtain all possible annotations for each sample (following Alg. 1). Moreover, we employ a passive learning (PL) approach as an upper bound by training the model on the full dataset, thus observing all available samples and annotations. Our baselines follow the analogous cross-validation setup.
4.3 Evaluation metrics
The ACAL strategies aim to guide the algorithm to model a representative distribution of the annotator’s perspectives while reducing human annotation effort. To this end, we evaluate the model both with a traditional evaluation metric and a metric aimed at comparing predicted and annotated distributions:
- Macro -score ()
For each sample in the test set, we select the label predicted by the model with the highest confidence, determine the golden label through a majority agreement aggregation, and compute the resulting macro -score.
- Jensen-Shannon Divergence ()
The measures the divergence between the distribution of label annotation and prediction (Nie et al., 2020). We report the average for the samples in the test set to measure how well the algorithm can model the annotation distribution.
Next, since our proposed annotator selection strategies aim to promote diversity, we introduce novel annotator-centric evaluation metrics. First, we report the average among annotators. Second, in line with Rawls’ principle of maximum fairness Rawls (1973), the result for the worst-off annotators:
- Per-annotator ()
We compute the for each annotator in the test set using their annotations as golden labels, and average it.
- Per-annotator ()
We compute the for each annotator in the test set using their annotations as target distribution, and average it.
- Worst per-annotator ()
We compute the for each annotator in the test set using their annotations as golden labels, and report the average of the lowest 10% (to mitigate noise).
- Worst per-annotator ()
We compute the for each annotator in the test set using their annotations as target distribution, and report the average of the lowest 10% (to mitigate noise).
These evaluation metrics allow us to measure the trade-offs between modeling the majority agreement, a representative distribution of annotations, and accounting for minority voices. We report these metrics on the validation set (as progress over the AL iterations) and test set (by using the best-performing model on the validation set), as described in Section 4.2.
5 Results
5.1 Test sets results
See Figure 2 for the performance of the DICES, MFTC, and MHS, respectively. For MFTC, we initially focus on care, since it is the task with neither the most nor least amount of data. For MHS, we start with dehumanize, since it saw the most medium-level disagreement. The rest of the results can be observed in Appendix B.
Combining our results across datasets, we see that data characteristics influence whether ACAL can learn performant models efficiently. In particular, we see that for DICES and MHS, ACAL may learn models that perform well using less data (38% and 62% reduction at best, respectively). Conversely, for MFTC, there is little impact of using ACAL over PL (5.6% less data used). A similar pattern holds when comparing ACAL to AL, though AL seems to be a strong baseline for MHS, where random sample selection leads to more efficient data usage (60%). AL with uncertainty sampling is more efficient for MFTC (13%).
When we compare the performance metrics, we see that the distributions obtained through ACAL are consistently closer to the ground truth distribution in DICES, as measured by than PL and AL. However, this pattern is not visible for MFTC and MHS. In terms of majority-voted , ACAL again leads to better scores in both DICES and MHS. Since DICES and MHS are datasets with moderate disagreement, we may benefit from using ACAL in such scenarios. Further, if the dataset contains a large number of annotators per sample, annotator selection strategies are shown to pick a more informative set of annotators to learn from.
We highlight some further dataset-specific findings that shed light on the differences between the annotator selection strategies in ACAL. First, in DICES, we see that for three out of four annotator sampling strategies (, , ), the choice of data sampling strategy has no impact on the performance of the model due to the low number of samples to choose from. Furthermore, only performs worse in terms of , overrepresenting outlier annotators. This hints that selecting annotators based on the average embedding of the annotated content strongly emphasizes diverging label behavior. Second, MFTC was annotated by a limited fixed set of annotators for whom we can construct a rich annotation history. However, since there are few annotators per sample to pick from, ACAL cannot leverage this information effectively. Again, we see that strategies perform relatively similarly to one another, except for the scores. Third, in MHS we observe that all strategies using random sample selection require less data. Since the task has low inter-rater agreement scores, uncertainty-based sampling wrongly attempts to sample annotations for correct high-entropy predictions, while this is an accurate distribution.
Our findings highlight that with many labels per sample, ACAL is more data-efficient than traditional AL and passive learning in terms of overall evaluation. However, when data characteristics differ and few annotations are available per sample, ACAL has less of an impact. Polarized agreement scores (either high agreement or no agreement) make the use of ACAL and AL cause little to no improvements over passive learning. This corroborates that (AC)AL leads to improvements in specific cases (Dor et al., 2020). Furthermore, we found conflicting results depending on the metric used ( and ). We closely examine the relationship between the evaluation metrics by turning to annotator-centric evaluation, observing how ACAL impacted predictions for individual annotators.
5.2 Annotator-centric evaluation
We show the annotator-centric evaluation metrics in Tables 2, 3, and 4 for DICES, MFTC (care) and MHS (dehumanize), respectively. We again describe per-dataset results. Again, for DICES and MHS, we observe a positive effect of using ACAL over PL, both in terms of data efficiency and final annotator-centric behavior. For these datasets, ACAL leads to a better representation of annotators on average (,), as well as a better representation of the 10% most different annotators (, ). Compared ACAL to AL, we mainly observe improvements in the DICES dataset, showing less data used and a better annotator-centric score. We observe a strong for the strategy and worse , corroborating our earlier finding that emphasizing diverging label behavior trades off with the averaged evaluation scores. Interestingly, this is not apparent in the scores. For MHS, all approaches using random data sampling () require considerably less data than passive learning. Further, since the pool of annotators for MHS is large (7K+), there will always be some annotators in disagreement with the output of our models, leading to a zero score on .
Average | Worst-off | ||||
App. | |||||
0.432 | 0.186 | 0.167 | 0.453 | -36.8 | |
0.424 | 0.187 | 0.155 | 0.450 | -32.7 | |
0.442 | 0.186 | 0.164 | 0.447 | -35.5 | |
0.431 | 0.203 | 0.169 | 0.370 | -30.0 | |
0.432 | 0.186 | 0.167 | 0.453 | -36.8 | |
0.424 | 0.187 | 0.155 | 0.450 | -32.7 | |
0.439 | 0.187 | 0.184 | 0.447 | -38.2 | |
0.431 | 0.203 | 0.169 | 0.370 | -30.0 | |
0.414 | 0.191 | 0.133 | 0.425 | -0.1 | |
0.384 | 0.192 | 0.117 | 0.427 | -0.1 | |
Passive | 0.371 | 0.211 | 0.123 | 0.479 | – |
Average | Worst-off | ||||
App. | |||||
0.611 | 0.141 | 0.377 | 0.247 | -1.6 | |
0.616 | 0.142 | 0.392 | 0.249 | -0.4 | |
0.600 | 0.145 | 0.351 | 0.248 | -1.7 | |
0.604 | 0.144 | 0.357 | 0.243 | -1.7 | |
0.612 | 0.143 | 0.377 | 0.252 | -5.6 | |
0.589 | 0.142 | 0.423 | 0.248 | -2.5 | |
0.608 | 0.143 | 0.399 | 0.258 | -1.1 | |
0.586 | 0.145 | 0.357 | 0.253 | -2.5 | |
0.586 | 0.141 | 0.392 | 0.255 | -0.2 | |
0.583 | 0.144 | 0.357 | 0.253 | -12.7 | |
Passive | 0.512 | 0.179 | 0.377 | 0.251 | – |
Average | Worst-off | ||||
App. | |||||
0.315 | 0.394 | 0.000 | 0.489 | -50.0 | |
0.322 | 0.397 | 0.000 | 0.478 | -62.5 | |
0.313 | 0.397 | 0.000 | 0.480 | -62.5 | |
0.318 | 0.398 | 0.000 | 0.479 | -62.5 | |
0.322 | 0.389 | 0.000 | 0.508 | -7.8 | |
0.328 | 0.388 | 0.000 | 0.507 | -7.8 | |
0.326 | 0.388 | 0.000 | 0.506 | -7.8 | |
0.326 | 0.384 | 0.000 | 0.513 | -3.0 | |
0.339 | 0.387 | 0.000 | 0.496 | -60.1 | |
0.331 | 0.390 | 0.000 | 0.497 | -24.7 | |
Passive | 0.202 | 0.424 | 0.000 | 0.547 | – |
5.3 Training plots
While the evaluation shows a pattern of efficient data use with ACAL under certain data conditions, it reveals little about how the metrics behave during training or how individual annotator strategies behave. To this end, we provide a complete overview of all metrics (as computed on the validation set) during training in App. B.3. Here we describe the major patterns reoccurring across our experiments using examples and show six of particular interest (Figure 3). Since the strategies only differ in what annotations are included during training, we only show plots related to the annotator-centric metrics.
We can see that there is an influence of both the data sampling and annotator strategies on the performance of the models. Only on DICES is the choice of irrelevant, probably due to the low number of samples. Specifically deteriorates slower for the worst-off annotators than the other strategies but does so without being able to uphold a competitive score. In MFTC, we see that when using , performance on dips at the start of training. Selecting annotators for samples with high predicted entropy initially leads to a decrease in average performance. The strategy seeks to first lower the entropy for the labels already encountered, though some of the variation in labels is irreconcilable. A similar reasoning holds for MHS, where the differences between strategies are even less impacted by the choice of . These two plots further underline our main finding that for ACAL to be impactful in representing diverse annotation perspectives, we need to ensure a (1) heterogeneous pool of annotators, and (2) a task that facilitates human label variation.
5.4 Change in task
In Fig. 4, we present a comparative analysis of two annotator-centric metrics across the three distinct tasks of MFTC and MHS, for which we have seen the least impact of ACAL over AL and PL. We cannot conclude that the chosen ACAL approach () offers a consistent improvement over sampling all annotations (), particularly given that the models using ACAL occasionally require more data to converge (Tables 8 to 11).
Initially, we hypothesized that tasks with a high degree of subjectivity would benefit from ACAL strategies, especially on metrics focused on the most marginalized (worst-off) annotators. These strategies typically involve selecting an annotator whose patterns of annotation diverge from the majority, either in terms of their annotation behavior or in the semantic content of their past annotations.
However, as depicted in Figure 4, when examining the task of dehumanize (high entropy), it becomes apparent that ACAL does not consistently outperform AL. ACAL demonstrates a lower -score than AL for this task, and on the other hand, a higher -score for a task that is less subjective, such as genocide. Similarly, when evaluating loyalty, which involves the moral dimension with the highest disagreement among annotators, the lower 10% of annotators are more accurately approximated with PL. This suggests that the effectiveness of annotation strategies varies depending on the task’s degree of subjectivity and available pool of annotators. The more heterogeneous the annotation behavior, indicative of a highly subjective task, the larger the pool of annotators required for each item selection. However, due to the limited annotations available per item in both datasets MFTC and MHS, even carefully selecting specific annotators may not adequately represent divergent annotation behavior in general, which challenges the generalization to unseen data. Finally, we can observe that there is a trade-off between modeling the majority of annotators equally, as reflected in the -score and prioritizing the minority viewpoint (). A better performance in one aspect does not necessarily guarantee superiority in the other.
6 Conclusion
We introduce Annotator-Centric Active Learning (ACAL), an active learning approach that incorporates annotator selection strategies aimed at capturing label variation among annotators. We experiment with tasks across three different datasets, each leading to different ACAL behaviors. One of these datasets, DICES, is the most realistic application of ACAL since the pool of possible annotators is the largest. Here, ACAL leads to more diverse label distributions using fewer annotations. However, we find that the effectiveness of the ACAL paradigm is contingent on data characteristics. These characteristics include the number of annotations per sample, the number of annotations per annotator, and the nature of disagreement in the task annotations. Our analysis shows that we can use these conditions to help explain the often disappointing results for AL in NLP applications.
Including annotator-centric evaluation reveals how methods with similar averaged performance deal with different levels of disagreement among annotators. We show that evaluation can be enhanced by focusing on individual annotators, as there is a large gap between conventional, averaged, and worst-off performance. Furthermore, many aspects of our ACAL approach can be experimented with, e.g. by swapping the order in which samples are selected (in our case first) and annotators (second), or investigating the impact of including annotator-specific demographic information, as it is inconsistently predictive of annotation behavior Orlikowski et al. (2023); Beck et al. (2024).
Limitations
The main limitation of this work is that the experiments are based on simulated active learning which is known to bear potential issues (Margatina and Aletras, 2023). In our study, a primary challenge arises with two of the datasets (MFTC, MHS), which, despite having a large pool of annotators, lack annotations from every annotator for each item. Consequently, in real-world scenarios, the annotator selection strategies for these datasets would benefit from access to a more extensive pool of annotators. This limitation likely contributes to the underperformance of ACAL on these datasets compared to DICES. We emphasize the need for more datasets that feature a greater number of annotations per item, as this would significantly enhance research efforts aimed at modeling human disagreement.
Since we evaluate four different annotator selection strategies and two sample selection strategies across three datasets and seven tasks, the amount of experiments is high. This did not allow for further investigation of the difference using different classification models, the extensive turning of hyperparameters, or even different training paradigms. Lastly, a limitation of our annotator selection strategies is that they rely on a small annotation history. This is why we require a warmup phase for some of the strategies, for which we decided to take a random sample of annotations. Incorporating more informed warmup strategies or incorporating ACAL strategies that do not rely on annotator history may positively impact its performance and data efficiency.
Ethical Considerations
Our goal is to approximate a good representation of human judgments over subjective tasks. We want to highlight the fact that the performance of the models differs a lot depending on which metric is used. We tried to account for a less majority-focussed view when evaluating the models which is very important, especially for more human-centered applications, such as hate-speech detection. However, the evaluation metrics we use do not fully capture the diversity of human judgments. The selection of metrics should align with the specific goals and motivations of the application, and there is a pressing need to develop more metrics to accurately reflect human variability in these tasks.
Our experiments are conducted on English datasets due to the scarcity of unaggregated datasets in other languages. In principle, ACAL can be applied to other languages (given the availability of multilingual models to semantically embed textual items for some particular strategies used in this work). We encourage the community to enrich the dataset landscape by incorporating more perspective-oriented datasets in various languages, ACAL potentially offers a more efficient method for creating such datasets in real-world scenarios.
Appendix A Detailed Experimental Setup
A.1 Cross validation details
We split the data on samples, meaning that all annotations for any given sample are completely contained in each separate split.
A.2 Hyperparameters
We report the hyperparameters for training passive, AL, and ACAL in Tables 5, 6, and 7, respectively. For turning the learning rate for passive learning, on each dataset, we started with a learning rate of 1e-06 and increased it by a factor of 3 in steps until the model showed a tendency to overfit quickly (within a single epoch). All other parameters are kept on their default setting.
Parameter | Value |
learning rate | 1e-04 (constant) |
max epochs | 50 |
early stopping | 3 |
batch size | 128 |
weight decay | 0.01 |
Parameter | Dataset (task) | Value |
learning rate | all | 1e-05 |
batch size | all | 128 |
epochs per round | all | 20 |
num rounds | all | 10 |
sample size | DICES | 79 |
sample size | MFTC (care) | 674 |
sample size | MFTC (betrayal) | 1011 |
sample size | MFTC (loyalty) | 263 |
sample size | MHS (dehumanize), MHS (genocide), MHS (respect) | 1728 |
Parameter | Dataset | Value |
learning rate | all | 1e-05 |
num rounds | DICES | 50 |
num rounds | MFTC (all), MHS (all) | 20 |
epochs per round | DICES, MHS (all) | 20 |
epochs per round | MFTC (all) | 30 |
sample size | DICES | 792 |
sample size | MFTC (care) | 1250 |
sample size | MFTC (betrayal) | 1894 |
sample size | MFTC (loyalty) | 512 |
sample size | MHS (dehumanize), MHS (genocide), MHS (respect) | 2899 |
A.3 Training details
Experiments were largely run between January and April 2024. Obtaining the ACAL results for a single run takes up to an hour on a Nvidia RTX4070. For large-scale computation, our experiments were run on a cluster with heterogeneous computing infrastructure, including RTX2080 Ti, A100, and Tesla T4 GPUs. Obtaining the results of all experiments required a total of 231 training runs, combining: (1) two data sampling strategies, (2) four annotator sampling strategies, plus an additional Oracle-based AL approach, (3) a passive learning approach. Each of the above were run for (1) three folds, each with a different seed, and (2) the seven tasks across three datasets. For training all our models, we employ the AdamW optimizer (Loshchilov and Hutter, 2018). Our code is based on the Huggingface library (Wolf et al., 2019), unmodified values are taken from their defaults.
A.4 ACAL Annotator Strategy details
Some of the strategies used for selecting annotators to provide a label to a sample
uses a sentence embedding model to represent the content that an annotator has annotated. We use all-MiniLM-L6-v2111https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. We select annotators that have not annotated yet (empty history) before picking from those with a history to prioritize filling the annotation history for each annotator.
creates an average embedding for the content annotated by each annotator and selects the most different annotator. We use the same sentence embedding model as . To avoid overfitting, we perform PCA and retain only the top 10 most informative principal components for representing each annotator.
A.5 Disagreement rates
We report the average disagreement rates per dataset and task in Figure 5, for each of the dataset and task combinations.
Appendix B Detailed Results Overview
B.1 Test set evaluation other MFTC and MHS tasks
See Table 6 for the trade-off between data efficiency and test-set performance for the two conventional metrics ( and ). We include copy the earlier mentioned results for MFTC (care) and MHS (dehumanize) for convenience.
B.2 Annotator-Centric evaluation for other MFTC and MHS tasks
We show the full annotator-centric metrics results for MFTC betrayal (Table 8), MFTC loyalty (Table 9), MHS genocide (Table 10), and MHS respect (Table 11).
Average | Worst-off | ||||
App. | |||||
0.578 | 0.147 | 0.420 | 0.199 | -1.6 | |
0.581 | 0.149 | 0.433 | 0.212 | -1.6 | |
0.593 | 0.161 | 0.430 | 0.239 | -5.0 | |
0.583 | 0.148 | 0.429 | 0.199 | -1.6 | |
0.594 | 0.150 | 0.419 | 0.203 | -2.5 | |
0.584 | 0.148 | 0.434 | 0.200 | -1.3 | |
0.588 | 0.149 | 0.435 | 0.204 | -1.0 | |
0.591 | 0.149 | 0.428 | 0.194 | -2.6 | |
0.589 | 0.147 | 0.431 | 0.195 | -48.6 | |
0.589 | 0.149 | 0.430 | 0.200 | -0.0 | |
passive | 0.481 | 0.199 | 0.360 | 0.290 | 0.0 |
Average | Worst-off | ||||
App. | |||||
0.564 | 0.177 | 0.222 | 0.372 | -0.4 | |
0.563 | 0.176 | 0.222 | 0.374 | -0.3 | |
0.573 | 0.176 | 0.222 | 0.370 | -0.3 | |
0.551 | 0.175 | 0.222 | 0.373 | -0.3 | |
0.557 | 0.177 | 0.217 | 0.357 | -1.1 | |
0.541 | 0.177 | 0.222 | 0.355 | -0.8 | |
0.556 | 0.177 | 0.222 | 0.358 | -0.9 | |
0.558 | 0.177 | 0.222 | 0.358 | -1.3 | |
0.560 | 0.176 | 0.222 | 0.361 | -29.1 | |
0.559 | 0.177 | 0.222 | 0.366 | -0.1 | |
passive | 0.512 | 0.183 | 0.261 | 0.309 | 0.0 |
Average | Worst-off | ||||
App. | |||||
0.700 | 0.227 | 0.000 | 0.560 | -6.3 | |
0.698 | 0.225 | 0.000 | 0.565 | -1.7 | |
0.700 | 0.224 | 0.000 | 0.566 | -1.7 | |
0.702 | 0.224 | 0.000 | 0.565 | -1.7 | |
0.711 | 0.229 | 0.000 | 0.549 | -12.6 | |
0.707 | 0.231 | 0.000 | 0.548 | -7.9 | |
0.709 | 0.231 | 0.000 | 0.548 | -7.9 | |
0.712 | 0.229 | 0.000 | 0.547 | -12.6 | |
0.339 | 0.387 | 0.000 | 0.496 | -60.1 | |
0.331 | 0.390 | 0.000 | 0.497 | -24.7 | |
passive | 0.700 | 0.245 | 0.000 | 0.570 | – |
Average | Worst-off | ||||
App. | |||||
0.460 | 0.331 | 0.000 | 0.528 | -18.8 | |
0.456 | 0.331 | 0.000 | 0.530 | -18.8 | |
0.461 | 0.331 | 0.000 | 0.529 | -18.8 | |
0.460 | 0.331 | 0.000 | 0.528 | -18.8 | |
0.466 | 0.323 | 0.000 | 0.533 | -4.9 | |
0.463 | 0.323 | 0.000 | 0.532 | -4.9 | |
0.459 | 0.324 | 0.000 | 0.531 | -4.9 | |
0.462 | 0.324 | 0.000 | 0.532 | -4.9 | |
0.339 | 0.387 | 0.000 | 0.496 | -60.1 | |
0.331 | 0.390 | 0.000 | 0.497 | -24.7 | |
passive | 0.259 | 0.405 | 0.000 | 0.587 | – |
B.3 Training process
In our main paper, we report a condensed version of all metrics during the training phase of the active learning approaches. Below, we provide a complete overview of all approaches over all metrics. The results can be seen in Figures 7 through 13.

### v2

@@
 Abstract
-To accurately capture the variability in human judgments for subjective NLP tasks, incorporating a wide range of perspectives in the annotation process is crucial.
 Active Learning (AL) addresses the high costs of collecting human annotations by strategically annotating the most informative samples.
+However, for subjective NLP tasks, incorporating a wide range of perspectives in the annotation process is crucial to capture the variability in human judgments.
 We introduce Annotator-Centric Active Learning (ACAL), which incorporates an annotator selection strategy following data sampling.
@@
 A challenging aspect of natural language understanding (NLU) is the variability of human judgment and interpretation in subjective tasks (e.g., hate speech detection) (Plank, 2022).
-While humans can navigate subjectivity naturally, most machine learning methods are insensitive to individual differences (Sandri et al., 2023) and underrepresented perspectives (van der Meer et al., 2024).
-Modern NLU approaches are commonly trained and tested on annotated datasets.
-In a subjective task, each data sample is typically labeled by a set of annotators, and differences in annotation are reconciled through aggregation techniques (e.g., majority voting), resulting in a single “gold label” (Uma et al., 2021).
-This approach, though effective for training ML algorithms, neglects the labels of minorities, which becomes problematic, especially, in the case of sensitive subjective tasks.
-Subjectivity has been addressed by modeling the full distribution of annotations for each data sample as opposed to aggregating them (Plank, 2022).
-However, resources for such approaches are scarce, as most datasets do not (yet) make fine-grained annotation details available (Cabitza et al., 2023), and representing a full range of perspectives is contingent on obtaining annotations from a diverse crowd (Bakker et al., 2022).
-One way of accounting for a limited annotation budget is to use Active Learning (Settles, 2012, AL).
-Given a pool of unannotated data samples, AL employs a sample selection strategy to select maximally informative samples for training, retrieving the corresponding annotations from a ground truth oracle (e.g., a single human expert).
-However, in subjective tasks there is no such oracle, instead we rely on a set of available annotators.
-Given this practical constraint, we argue that informativeness for AL manifests in both samples and annotations, as the model should also be guided to reflect the distribution of annotations.
-Demanding all available annotators to annotate all selected samples would provide a truthful representation of the annotation distribution, but is often unfeasible, especially if the pool of annotators is large.
-Thus, deciding which annotator(s) should annotate the selected samples is as critical as selecting which samples to annotate.
-We introduce Annotator-Centric Active Learning (ACAL) to account for annotation diversity in subjective tasks.
-In ACAL, the sample selection strategy of traditional AL is followed by an annotator selection strategy as Figure 1 shows.
-For each data sample selected through the sample selection strategy, the annotator selection strategy selects an annotator from the available annotators.
-We make the following contributions: (1) We create ACAL, extending the AL approach to optimize for diversity among annotators when learning soft labels in subjective tasks.
-(2) We introduce a suite of annotator-centric evaluation metrics to measure both representativeness and diversity.
-(3) We demonstrate our approach’s effectiveness on three diverse datasets with subjective tasks—hate speech detection, moral value classification, and safety judgments.
-Our experiments show that ACAL works to better approximate the distribution of human judgments with a lower annotation budget.
-However, this effectiveness requires a large pool of diverse annotators, as is the case for one of our datasets.
-In other cases, the differences between ACAL and traditional AL become smaller.
-Through our annotator-centric evaluation, we show that task agreement and the number of available annotations both influence the effectiveness of ACAL, hinting at a direct trade-off between learning from a majority versus being sensitive to minority annotations.
+In a subjective task, a data sample is typically labeled by a set of annotators, and differences in annotation are reconciled via majority voting, resulting in a single (supposedly, true) “gold label” (Uma et al., 2021).
+However, this approach has been criticized for treating label variation exclusively as noise, which is especially problematic in sensitive subjective tasks (Aroyo and Welty, 2015) since it can lead to exclusion of minority voices (Leonardelli et al., 2021).
+Subjectivity can be addressed by modeling the full distribution of annotations for each data sample instead of employing gold labels (Plank, 2022).
+However, resources for such approaches are scarce, as most datasets do not (yet) make fine-grained annotation details available (Cabitza et al., 2023), and representing a full range of perspectives is contingent on obtaining costly annotations from a diverse set of annotators (Bakker et al., 2022).
+One way to handle a limited annotation budget is to use Active Learning (Settles, 2012, AL).
+Given a pool of unannotated data samples, AL employs a sample selection strategy to obtain maximally informative samples, retrieving the corresponding annotations from a ground truth oracle (e.g., a single human expert).
+However, in subjective tasks, there is no such oracle.
+Instead, we rely on a set of available annotators.
+Demanding all available annotators to annotate all samples would provide a truthful representation of the annotation distribution, but is often unfeasible, especially if the pool of annotators is large.
+Thus, deciding which annotator(s) should annotate is as critical as deciding which samples to annotate.
+In most practical applications, annotators are randomly selected.
+This results in an annotation distribution insensitive to outlier annotators —most annotations reflect the majority voices and fewer reflect the minority voices.
+This may not be desirable in applications such as hate speech, where the opinion of majority and minority should be valued equally.
+In such cases, a more deliberate annotator selection is required.
+To ensure a balanced representation of majority and minority voices, we can leverage strategies inspired by Rawls’ principle of fairness Rawls (1973), which advocates that a fair society is achieved when the well-being of the worst-off members of society (the minority annotators, in this case) is maximized.
+We introduce Annotator-Centric Active Learning (ACAL) to emphasize and control who annotates which sample.
+In ACAL (Figure 1), the sample selection strategy of traditional AL is followed by an annotator selection strategy, indicating which of the available annotators should annotate each selected data sample.
+Contributions
+Our experiments show that the proposed ACAL methods can approximate the distribution of human judgments similar to AL while requiring a lower annotation budget and modeling individual and minority voices more accurately.
+However, our evaluation shows how the task’s annotator agreement and the number of available annotations impact ACAL’s effectiveness—ACAL is most effective when a large pool of diverse annotators is available.
+Importantly, our experiments show how the ACAL framework controls how models learn to represent majority and minority annotations, which is crucial for subjective and sensitive applications.
 2 Related work
-We review related works on annotator disagreement and active learning.
-Our work is novel in combining these fields to (1) represent annotation distributions through soft labels, (2) incorporate annotator selection strategies in the active learning loop, and (3) evaluate with annotator-centric metrics next to traditional evaluation.
 2.1 Learning with annotator disagreement
 Modeling annotator disagreement is garnering increasing attention (Aroyo and Welty, 2015; Uma et al., 2021; Plank, 2022; Cabitza et al., 2023).
-For instance, some aggregation methods can lead to a fairer representation than simple majority (Hovy et al., 2013; Tao et al., 2018).
-Alternatively, the full annotation distribution can be modeled using soft labels (e.g., Peterson et al., 2019; Müller et al., 2019; Fornaciari et al., 2021; Collins et al., 2022).
-Other approaches leverage annotator-specific information, e.g., by including individual classification heads per annotator (Davani et al., 2022), embedding annotator-specific behavior (Mokhberian et al., 2023), or encoding the annotator’s socio-demographic information (Beck et al., 2023).
-Yet, representing annotator diversity remains challenging.
+Changing annotation aggregation methods can lead to a fairer representation than simple majority (Hovy et al., 2013; Tao et al., 2018).
+Alternatively, the full annotation distribution can be modeled using soft labels (Peterson et al., 2019; Müller et al., 2019; Collins et al., 2022).
+Other approaches leverage annotator-specific information, e.g., by including individual classification heads per annotator (Davani et al., 2022), embedding annotator behavior (Mokhberian et al., 2023), or encoding the annotator’s socio-demographic information (Beck et al., 2023).
+Representing annotator diversity remains challenging.
 Standard calibration metrics under human label variation may be unsuitable, especially when the variation is high (Baan et al., 2022).
@@
 (2024).
-Further, solely measuring differences among sociodemographic traits is not sufficient to fully capture opinion diversity (Orlikowski et al., 2023).
-To this end, we represent diversity based on which annotators have annotated, what they annotated, and how they have annotated.
-We experiment with different annotator selection strategies to reveal what aspects impact task performance and annotation budget.
+Further, solely measuring differences among sociodemographic traits is not sufficient to capture opinion diversity (Orlikowski et al., 2023).
+We represent diversity based on which annotators annotated what and how. We experiment with annotator selection strategies to reveal what aspects impact task performance and annotation budget.
 2.2 Active Learning
-AL enables a supervised learning model to achieve high performance with a few training examples if chosen judiciously Settles (2012).
-In a typical AL scenario, a vast collection of unlabeled data is available, and an oracle (e.g., a human expert) can be asked to annotate this unlabeled data.
-A sampling strategy is employed to iteratively (and smartly) select the next batch of unlabeled data for annotation by the oracle Ren et al.
+AL enables a supervised learning model to achieve high performance by judiciously choosing a few training examples Settles (2012).
+In a typical AL scenario, a large collection of unlabeled data is available, and an oracle (e.g., a human expert) is asked to annotate this unlabeled data.
+A sampling strategy is used to iteratively select the next batch of unlabeled data for annotation Ren et al.
 (2021).
-AL has found widespread application in the field of NLP Zhang et al.
+AL has found widespread application in NLP Zhang et al.
 (2022).
 Two main strategies are employed, either by selecting the unlabeled samples on which the model prediction is most uncertain (Zhang et al., 2017), or by selecting samples that are most representative of the unlabeled dataset (Erdmann et al., 2019; Zhao et al., 2020).
-The combination of AL and annotator diversity is a novel direction that has not garnered much attention yet.
-Existing work proposes to align model uncertainty with annotator uncertainty (Baumler et al., 2023), whereas others adapt annotator-specific classification heads in AL settings (Wang and Plank, 2023), or select texts to annotate based on annotator preferences (Kanclerz et al., 2023).
-Existing methods ignore a crucial part of learning with human variation: the diversity among annotators.
-We focus on which annotators should annotate, such that it best informs us about the underlying label diversity.
+The combination of AL and annotator diversity is a novel direction.
+Existing works propose to align model and annotator uncertainties (Baumler et al., 2023), adapt annotator-specific classification heads in AL settings (Wang and Plank, 2023), or select texts to annotate based on annotator preferences (Kanclerz et al., 2023).
+These methods ignore a crucial part of learning with human variation: the diversity among annotators.
+We focus on selecting annotators such that they best inform us about the underlying label diversity.
 3 Method
@@
 3.1 Soft-label prediction
-Consider a dataset composed of triples , where is a data sample (i.e., a piece of text) and is the class label assigned by annotator .
+Consider a dataset of triples , where is a data sample (i.e., a piece of text) and is the class label assigned by annotator .
 The multiple labels assigned to a sample by the different annotators are usually combined into an aggregated label .
@@
 1 provides an overview of the procedure.
-In our sample selection strategies, a batch of data of a given size is queried at each iteration.
-In our experiments, we compare the following strategies:
+In the sample selection strategies, a batch of data of a given size is queried at each iteration. Our experiments compare the following strategies:
 - Random ()
@@
 - Uncertainty ()
-predicts a distribution over class labels with for each . Select the samples with the highest prediction entropy (i.e., the samples on which the model is most uncertain).
+predicts a distribution over class labels with for each , and selects samples with the highest prediction entropy (the samples the model is most uncertain about).
 3.3 Annotator-Centric Active Learning
-The ACAL approach builds on the AL approach.
+ACAL builds on AL.
 In contrast to AL, which retrieves an aggregated annotation , ACAL employs an annotator selection strategy to select one annotator and their annotation for each selected data point .
@@
 2 describes the ACAL approach.
-Dataset | Task (dimension) | Num. Samples | Num. Annotators | Num. Annotations | Avg. Annotations per item |
-DICES | Safety Judgment | 990 | 172 | 72,103 | 72.83 |
-MFTC | Morality (care) | 8434 | 23 | 31310 | 3.71 |
-MFTC | Morality (loyalty) | 3288 | 23 | 12803 | 3.89 |
-MFTC | Morality (betrayal) | 12546 | 23 | 47002 | 3.75 |
-MHS | Hate Speech (dehumanize, genocide, respect) | 17282 | 7807 | 57980 | 3.35 |
-We propose annotator selection strategies that include annotations from diverse annotators.
-The strategies vary in the type of information used to represent differences between annotators, and include what or how the annotators have annotated thus far.
-We test the following strategies:
+We propose three annotator selection strategies to gather a distribution that uniformly contains all possible (majority and minority) labels, inspired by Rawls’ principle of fairness Rawls (1973).
+The strategies vary in the type of information used to represent differences between annotators, including what or how the annotators have annotated thus far.
+Our experiments compare the following strategies:
 - Random ()
-selects one random annotator .
+randomly selects an annotator .
 - Label Minority ()
 considers only the labels that annotators have assigned.
+The minority label is selected as the class with the smallest annotation count in the available dataset thus far.
 Given a new sample , selects the available annotator that has the largest bias toward the minority label compared to the other available annotators, i.e., who has annotated other samples with the minority label the most.
-The minority label is selected as the class with the smallest annotation count in the available dataset thus far.
 - Semantic Diversity ()
@@
 4.1 Datasets
-Table 1 introduces the three datasets that we use, with variation in domain, annotation task (in italics), annotator count, and annotations per instance.
+We use three datasets which vary in domain, annotation task (in italics), annotator count, and annotations per instance.
 The DICES Corpus (Aroyo et al., 2023) is composed of 990 conversations with an LLM where 172 annotators provided judgments on whether a generated response can be deemed safe (3-way judgments: yes, no, unsure).
+Samples have 73 annotations on average.
 We perform a multi-class classification with the scores.
@@
 (2013).
-We select the elements of loyalty (lowest annotation count), care (medium count), and betrayal (highest count) and perform three binary classifications to predict the presence of the respective elements.
-As most tweets were labeled non-moral (i.e., with no moral element), we balanced the datasets by subsampling the non-moral class.
+We select the elements of loyalty (lowest annotation count), care (average count), and betrayal (highest count).
+Samples have 4 annotations on average.
+We create three binary classifications to predict the presence of the respective elements.
+As most tweets were labeled as non-moral (i.e., with no moral element), we balanced the datasets by subsampling the non-moral class.
 The MHS Corpus (Sachdeva et al., 2022) consists of 50K social media comments on which 8K annotators judged three hate speech aspects—dehumanize (low inter-rater agreement), respect (medium agreement), and genocide (high agreement)—on a 5-point Likert scale.
+Samples have 3 annotations on average.
 We perform a multi-class classification with the annotated Likert scores for each task.
-The datasets and tasks differ in the entropy scores over annotations (Appendix A.5).
-DICES and MHS generally have medium normalized entropy scores (most lie between , whereas the MFTC entropy scores are highly polarized.
-4.2 Training procedure
-We test the annotator selection strategies proposed in Section 3.3 by comparing all possible combinations of the two different sample selection strategies ( and ) with the annotator selection strategies (, , , and ).
-At each round, we use to select unique samples from the unlabeled data pool .
-We empirically select to be the smallest between 5% of the number of available annotations and the number of unique samples in the training set.
-For each selected sample , we use to select one annotator and retrieve their annotation .
-To populate the annotation history for the annotation selection strategies, we perform a single warmup round with randomly selected annotations before starting the ACAL iterations (Zhang et al., 2022).
-We report our training progress results on a validation set with 3-fold cross-validation, showing the average to account for stability across random data splits (into 80% train, 10% validation, and 10% test) and initialization.
-Then, we select the model iteration that led to the best performance (according to ) on the validation set and evaluate it using a separate test set.
-We compare our work with traditional Oracle-based AL approaches ( and ), which use the data sampling strategies but obtain all possible annotations for each sample (following Alg.
-1).
-Moreover, we employ a passive learning (PL) approach as an upper bound by training the model on the full dataset, thus observing all available samples and annotations.
-Our baselines follow the analogous cross-validation setup.
+The datasets and tasks differ in levels of annotator agreement, measured via entropy of the annotation distribution.
+DICES and MHS generally have medium entropy scores, whereas the MFTC entropy is highly polarized (divided between samples with very high and very low agreement).
+Appendix A.5 provides details of the entropy scores.
 4.2 Evaluation metrics
-The ACAL strategies aim to guide the algorithm to model a representative distribution of the annotator’s perspectives while reducing human annotation effort.
+The ACAL strategies aim to guide the model to learn a representative distribution of the annotator’s perspectives while reducing annotation effort.
 To this end, we evaluate the model both with a traditional evaluation metric and a metric aimed at comparing predicted and annotated distributions:
@@
 The measures the divergence between the distribution of label annotation and prediction (Nie et al., 2020).
-We report the average for the samples in the test set to measure how well the algorithm can model the annotation distribution.
-Next, since our proposed annotator selection strategies aim to promote diversity, we introduce novel annotator-centric evaluation metrics.
+We report the average for the samples in the test set to measure how well the model can represent the annotation distribution.
+Further, since ACAL shifts the focus to annotators, we introduce novel annotator-centric evaluation metrics.
 First, we report the average among annotators.
-Second, in line with Rawls’ principle of maximum fairness Rawls (1973), the result for the worst-off annotators:
-- Per-annotator ()
-We compute the for each annotator in the test set using their annotations as golden labels, and average it.
-- Per-annotator ()
-We compute the for each annotator in the test set using their annotations as target distribution, and average it.
-- Worst per-annotator ()
-We compute the for each annotator in the test set using their annotations as golden labels, and report the average of the lowest 10% (to mitigate noise).
-- Worst per-annotator ()
-We compute the for each annotator in the test set using their annotations as target distribution, and report the average of the lowest 10% (to mitigate noise).
-These evaluation metrics allow us to measure the trade-offs between modeling the majority agreement, a representative distribution of annotations, and accounting for minority voices.
-We report these metrics on the validation set (as progress over the AL iterations) and test set (by using the best-performing model on the validation set), as described in Section 4.2.
+Second, in line with Rawls’ principle of fairness, the result for the worst-off annotators:
+- Per-annotator () and ()
+We compute the (or ) for each annotator in the test set using their annotations as golden labels (or target distribution), and average it.
+- Worst per-annotator () and ()
+We compute the (or ) for each annotator in the test set using their annotations as golden labels (or target distribution), and report the average of the lowest 10% (to mitigate noise).
+These metrics allow us to measure the trade-offs between modeling the majority agreement, a representative distribution of annotations, and accounting for minority voices.
+In the next section, we describe how we obtained the results.
+4.3 Training procedure
+We test the annotator selection strategies proposed in Section 3.3 by comparing all combinations of the two sample selection strategies ( and ) and the four annotator selection strategies (, , , and ).
+At each iteration, we use to select unique samples from the unlabeled data pool .
+We select as the smallest between 5% of the number of available annotations and the number of unique samples in the training set.
+For each selected sample , we use to select one annotator and retrieve their annotation .
+We split each dataset into 80% train, 10% validation, and 10% test.
+We start the training procedure with a warmup iteration of randomly selected annotations (Zhang et al., 2022).
+We proceed with the ACAL iterations by combining and .
+We select the model iteration that led to the best performance on the validation set and evaluate it on the test set.
+We repeat this process across three data splits and model initializations.
+We report the average scores on the test set.
+Appendix A contains additional details on training.
+We compare ACAL with traditional oracle-based AL approaches ( and ), which use the data sampling strategies but obtain all possible annotations for each sample as in Alg.
+1.
+Further, we employ a passive learning (PL) approach as an upper bound by training the model on the full dataset, thus observing all available samples and annotations.
+Similar to ACAL, the AL and PL baselines are averaged over three seeds.
 5 Results
-5.1 Test sets results
-See Figure 2 for the performance of the DICES, MFTC, and MHS, respectively.
-For MFTC, we initially focus on care, since it is the task with neither the most nor least amount of data.
-For MHS, we start with dehumanize, since it saw the most medium-level disagreement.
-The rest of the results can be observed in Appendix B.
-Combining our results across datasets, we see that data characteristics influence whether ACAL can learn performant models efficiently.
-In particular, we see that for DICES and MHS, ACAL may learn models that perform well using less data (38% and 62% reduction at best, respectively).
-Conversely, for MFTC, there is little impact of using ACAL over PL (5.6% less data used).
-A similar pattern holds when comparing ACAL to AL, though AL seems to be a strong baseline for MHS, where random sample selection leads to more efficient data usage (60%).
-AL with uncertainty sampling is more efficient for MFTC (13%).
-When we compare the performance metrics, we see that the distributions obtained through ACAL are consistently closer to the ground truth distribution in DICES, as measured by than PL and AL.
-However, this pattern is not visible for MFTC and MHS.
-In terms of majority-voted , ACAL again leads to better scores in both DICES and MHS.
-Since DICES and MHS are datasets with moderate disagreement, we may benefit from using ACAL in such scenarios.
-Further, if the dataset contains a large number of annotators per sample, annotator selection strategies are shown to pick a more informative set of annotators to learn from.
-We highlight some further dataset-specific findings that shed light on the differences between the annotator selection strategies in ACAL.
-First, in DICES, we see that for three out of four annotator sampling strategies (, , ), the choice of data sampling strategy has no impact on the performance of the model due to the low number of samples to choose from.
-Furthermore, only performs worse in terms of , overrepresenting outlier annotators.
-This hints that selecting annotators based on the average embedding of the annotated content strongly emphasizes diverging label behavior.
-Second, MFTC was annotated by a limited fixed set of annotators for whom we can construct a rich annotation history.
-However, since there are few annotators per sample to pick from, ACAL cannot leverage this information effectively.
-Again, we see that strategies perform relatively similarly to one another, except for the scores.
-Third, in MHS we observe that all strategies using random sample selection require less data.
-Since the task has low inter-rater agreement scores, uncertainty-based sampling wrongly attempts to sample annotations for correct high-entropy predictions, while this is an accurate distribution.
-Our findings highlight that with many labels per sample, ACAL is more data-efficient than traditional AL and passive learning in terms of overall evaluation.
-However, when data characteristics differ and few annotations are available per sample, ACAL has less of an impact.
-Polarized agreement scores (either high agreement or no agreement) make the use of ACAL and AL cause little to no improvements over passive learning.
-This corroborates that (AC)AL leads to improvements in specific cases (Dor et al., 2020).
-Furthermore, we found conflicting results depending on the metric used ( and ).
-We closely examine the relationship between the evaluation metrics by turning to annotator-centric evaluation, observing how ACAL impacted predictions for individual annotators.
-5.2 Annotator-centric evaluation
-We show the annotator-centric evaluation metrics in Tables 2, 3, and 4 for DICES, MFTC (care) and MHS (dehumanize), respectively.
-We again describe per-dataset results.
-Again, for DICES and MHS, we observe a positive effect of using ACAL over PL, both in terms of data efficiency and final annotator-centric behavior.
-For these datasets, ACAL leads to a better representation of annotators on average (,), as well as a better representation of the 10% most different annotators (, ).
-Compared ACAL to AL, we mainly observe improvements in the DICES dataset, showing less data used and a better annotator-centric score.
-We observe a strong for the strategy and worse , corroborating our earlier finding that emphasizing diverging label behavior trades off with the averaged evaluation scores.
-Interestingly, this is not apparent in the scores.
-For MHS, all approaches using random data sampling () require considerably less data than passive learning.
-Further, since the pool of annotators for MHS is large (7K+), there will always be some annotators in disagreement with the output of our models, leading to a zero score on .
-Average | Worst-off | ||||
-App. | |||||
-0.432 | 0.186 | 0.167 | 0.453 | -36.8 | |
-0.424 | 0.187 | 0.155 | 0.450 | -32.7 | |
-0.442 | 0.186 | 0.164 | 0.447 | -35.5 | |
-0.431 | 0.203 | 0.169 | 0.370 | -30.0 | |
-0.432 | 0.186 | 0.167 | 0.453 | -36.8 | |
-0.424 | 0.187 | 0.155 | 0.450 | -32.7 | |
-0.439 | 0.187 | 0.184 | 0.447 | -38.2 | |
-0.431 | 0.203 | 0.169 | 0.370 | -30.0 | |
-0.414 | 0.191 | 0.133 | 0.425 | -0.1 | |
-0.384 | 0.192 | 0.117 | 0.427 | -0.1 | |
-Passive | 0.371 | 0.211 | 0.123 | 0.479 | – |
-Average | Worst-off | ||||
-App. | |||||
-0.611 | 0.141 | 0.377 | 0.247 | -1.6 | |
-0.616 | 0.142 | 0.392 | 0.249 | -0.4 | |
-0.600 | 0.145 | 0.351 | 0.248 | -1.7 | |
-0.604 | 0.144 | 0.357 | 0.243 | -1.7 | |
-0.612 | 0.143 | 0.377 | 0.252 | -5.6 | |
-0.589 | 0.142 | 0.423 | 0.248 | -2.5 | |
-0.608 | 0.143 | 0.399 | 0.258 | -1.1 | |
-0.586 | 0.145 | 0.357 | 0.253 | -2.5 | |
-0.586 | 0.141 | 0.392 | 0.255 | -0.2 | |
-0.583 | 0.144 | 0.357 | 0.253 | -12.7 | |
-Passive | 0.512 | 0.179 | 0.377 | 0.251 | – |
-Average | Worst-off | ||||
-App. | |||||
-0.315 | 0.394 | 0.000 | 0.489 | -50.0 | |
-0.322 | 0.397 | 0.000 | 0.478 | -62.5 | |
-0.313 | 0.397 | 0.000 | 0.480 | -62.5 | |
-0.318 | 0.398 | 0.000 | 0.479 | -62.5 | |
-0.322 | 0.389 | 0.000 | 0.508 | -7.8 | |
-0.328 | 0.388 | 0.000 | 0.507 | -7.8 | |
-0.326 | 0.388 | 0.000 | 0.506 | -7.8 | |
-0.326 | 0.384 | 0.000 | 0.513 | -3.0 | |
-0.339 | 0.387 | 0.000 | 0.496 | -60.1 | |
-0.331 | 0.390 | 0.000 | 0.497 | -24.7 | |
-Passive | 0.202 | 0.424 | 0.000 | 0.547 | – |
-5.3 Training plots
-While the evaluation shows a pattern of efficient data use with ACAL under certain data conditions, it reveals little about how the metrics behave during training or how individual annotator strategies behave.
-To this end, we provide a complete overview of all metrics (as computed on the validation set) during training in App.
-B.3.
-Here we describe the major patterns reoccurring across our experiments using examples and show six of particular interest (Figure 3).
-Since the strategies only differ in what annotations are included during training, we only show plots related to the annotator-centric metrics.
-We can see that there is an influence of both the data sampling and annotator strategies on the performance of the models.
-Only on DICES is the choice of irrelevant, probably due to the low number of samples.
-Specifically deteriorates slower for the worst-off annotators than the other strategies but does so without being able to uphold a competitive score.
-In MFTC, we see that when using , performance on dips at the start of training.
-Selecting annotators for samples with high predicted entropy initially leads to a decrease in average performance.
-The strategy seeks to first lower the entropy for the labels already encountered, though some of the variation in labels is irreconcilable.
-A similar reasoning holds for MHS, where the differences between strategies are even less impacted by the choice of .
-These two plots further underline our main finding that for ACAL to be impactful in representing diverse annotation perspectives, we need to ensure a (1) heterogeneous pool of annotators, and (2) a task that facilitates human label variation.
-5.4 Change in task
-In Fig.
-4, we present a comparative analysis of two annotator-centric metrics across the three distinct tasks of MFTC and MHS, for which we have seen the least impact of ACAL over AL and PL.
-We cannot conclude that the chosen ACAL approach () offers a consistent improvement over sampling all annotations (), particularly given that the models using ACAL occasionally require more data to converge (Tables 8 to 11).
-Initially, we hypothesized that tasks with a high degree of subjectivity would benefit from ACAL strategies, especially on metrics focused on the most marginalized (worst-off) annotators.
-These strategies typically involve selecting an annotator whose patterns of annotation diverge from the majority, either in terms of their annotation behavior or in the semantic content of their past annotations.
-However, as depicted in Figure 4, when examining the task of dehumanize (high entropy), it becomes apparent that ACAL does not consistently outperform AL.
-ACAL demonstrates a lower -score than AL for this task, and on the other hand, a higher -score for a task that is less subjective, such as genocide.
-Similarly, when evaluating loyalty, which involves the moral dimension with the highest disagreement among annotators, the lower 10% of annotators are more accurately approximated with PL.
-This suggests that the effectiveness of annotation strategies varies depending on the task’s degree of subjectivity and available pool of annotators.
-The more heterogeneous the annotation behavior, indicative of a highly subjective task, the larger the pool of annotators required for each item selection.
-However, due to the limited annotations available per item in both datasets MFTC and MHS, even carefully selecting specific annotators may not adequately represent divergent annotation behavior in general, which challenges the generalization to unseen data.
-Finally, we can observe that there is a trade-off between modeling the majority of annotators equally, as reflected in the -score and prioritizing the minority viewpoint ().
-A better performance in one aspect does not necessarily guarantee superiority in the other.
+We start by highlighting the benefits of ACAL over AL and PL (Section 5.1).
+Next, we closely examine ACAL on efficiency and fairness (Section 5.2).
+Then, we select a few cases of interest and dive deeper into the strategies’ behavior during training (Section 5.3).
+Finally, we investigate ACAL across varying levels of subjectivity (Section 5.4).
+[width=.5]tikz/snapshot
+5.1 Highlights
+Our experiments show that ACAL can have a beneficial impact over using PL and AL.
+Figure 2 highlights two main findings: (1) ACAL strategies can more quickly learn to represent the annotation distribution with a large pool of annotators, and (2) when agreement between annotators is polarized, ACAL leads to improved results compared to learning from aggregated labels.
+In the next sections, we provide a deeper understanding of the conditions in which ACAL works well.
+5.2 Efficiency and Fairness
+Table 1 presents the results of evaluating the best models on the test set.
+We analyze the results along two dimensions: (a) efficiency: what is the impact of the different strategies on the trade-off between annotation budget and performance?
+(b) fairness: do the selection strategies that aim for a balanced consideration of minority and majority views lead to better performance in the human-centric evaluation metrics?
+For MFTC we focus on care because it has an average number of samples available, and for MHS we focus on dehumanize because it has high levels of disagreement.
+Appendix B presents additional results.
+Average | Worst-off | |||||||
+App. | ||||||||
+DICES | 53.2 | .100 | 43.2 | .186 | 16.7 | .453 | -36.8 | |
+55.5 | .101 | 42.4 | .187 | 15.5 | .450 | -32.7 | ||
+61.0 | .103 | 44.2 | .186 | 16.4 | .447 | -35.5 | ||
+58.9 | .142 | 43.1 | .203 | 16.9 | .370 | -30.0 | ||
+53.2 | .100 | 43.2 | .186 | 16.7 | .453 | -36.8 | ||
+55.5 | .101 | 42.4 | .187 | 15.5 | .450 | -32.7 | ||
+63.1 | .098 | 43.9 | .187 | 18.4 | .447 | -38.2 | ||
+58.9 | .142 | 43.1 | .203 | 16.9 | .370 | -30.0 | ||
+59.1 | .112 | 41.4 | .191 | 13.3 | .425 | -0.1 | ||
+46.2 | .110 | 38.4 | .192 | 11.7 | .427 | -0.1 | ||
+PL | 59.0 | .105 | 37.1 | .211 | 12.3 | .479 | – | |
+MFTC (care) | 78.9 | .038 | 61.1 | .141 | 37.7 | .247 | -1.6 | |
+78.5 | .037 | 61.6 | .142 | 39.2 | .249 | -0.4 | ||
+78.1 | .039 | 60.0 | .145 | 35.1 | .248 | -1.7 | ||
+76.6 | .040 | 60.4 | .144 | 35.7 | .243 | -1.7 | ||
+79.4 | .038 | 61.2 | .143 | 37.7 | .252 | -5.6 | ||
+80.7 | .037 | 58.9 | .142 | 42.3 | .248 | -2.5 | ||
+79.1 | .037 | 60.8 | .143 | 39.9 | .258 | -1.1 | ||
+78.1 | .040 | 58.6 | .145 | 35.7 | .253 | -2.5 | ||
+79.0 | .037 | 58.6 | .141 | 39.2 | .255 | -0.2 | ||
+79.4 | .037 | 58.3 | .144 | 35.7 | .253 | -12.7 | ||
+PL | 81.1 | .032 | 51.2 | .179 | 37.7 | .251 | – | |
+MHS (dehumanize) | 33.6 | .081 | 31.5 | .394 | 0.0 | .489 | -50.0 | |
+33.1 | .081 | 32.2 | .397 | 0.0 | .478 | -62.5 | ||
+30.5 | .079 | 31.3 | .397 | 0.0 | .480 | -62.5 | ||
+32.4 | .081 | 31.8 | .398 | 0.0 | .479 | -62.5 | ||
+32.4 | .080 | 32.2 | .389 | 0.0 | .508 | -7.8 | ||
+33.1 | .080 | 32.8 | .388 | 0.0 | .507 | -7.8 | ||
+33.6 | .080 | 32.6 | .388 | 0.0 | .506 | -7.8 | ||
+33.0 | .079 | 32.6 | .384 | 0.0 | .513 | -3.0 | ||
+32.8 | .077 | 33.9 | .387 | 0.0 | .496 | -60.1 | ||
+33.3 | .080 | 33.1 | .390 | 0.0 | .497 | -24.7 | ||
+PL | 28.0 | .075 | 20.2 | .424 | 0.0 | .547 | – |
+Efficiency
+We discuss the performance on and to measure how well the proposed strategies model label distributions and examine the used annotator budget.
+Across all tasks and datasets, ACAL and AL consistently yield comparable or superior and with a lower annotation budget than PL.
+When comparing ACAL with AL, the results vary depending on the task and dataset.
+For DICES, there is a significant benefit to using ACAL, as it can save up to 40% of the annotation budget while yielding better scores across all metrics than AL.
+With AL, we observe only a small reduction in annotation cost.
+For MFTC, AL with leads to the largest cost benefits (12% less annotation budget), but at a cost in terms of absolute and .
+ACAL slightly outperforms AL but does not lead to a decrease in annotation budget.
+For MHS, both AL and ACAL significantly reduce the annotation cost (60%) while yielding better scores than PL—however, AL and ACAL do not show substantial performance differences.
+Overall, we conclude that ACAL is most efficient when the pool of available annotators for one sample is large (as with the DICES dataset), whereas the difference between ACAL and AL is negligible with a small pool of annotators per data sample (as with MFTC and MHS).
+Fairness
+We investigate the extent to which the models represent individual annotators fairly and capture minority opinions via the annotator-centric evaluation metrics (, , , and ).
+We observe a substantial improvement when using AL or ACAL over PL.
+Further, we observe no single winner-takes-all approach: high and scores do not consistently cooccur with high scores for the annotator-centric metrics.
+This highlights the need for a more comprehensive evaluation to assess models for subjective tasks.
+We observe that ACAL slightly outperforms AL in modeling individual annotators ( and ).
+This trend is particularly evident with DICES, again likely due to the large pool of annotators available per data sample.
+Lastly, ACAL is best in the worst-off metrics ( and ), showing the ability to better represent minority opinions as a direct consequence of the proposed annotator selection strategies on DICES and MFTC.
+However, all approaches score 0 for on MHS.
+This is due to the high disagreement in this dataset: the 10% worst-off annotators always disagree with a hard label derived from the predicted label distribution.
+In conclusion, our experiments show that, when a large pool of annotators is available, a targeted sampling of annotators requires fewer annotations and is fairer.
+That is, minority opinions are better represented without large sacrifices in performance compared to the overall label distribution.
+[width=]tikz/training_process_selected
+5.3 Convergence
+The evaluation on the test set paints a general picture of the advantage of using ACAL over AL or PL.
+In this section, we assess how different ACAL strategies converge over iterations.
+We describe the major patterns across our experiments by analyzing six examples of interest with and (Figure 3).
+We select because it reveals how well individual annotators are modeled on average, and to measure how strategies deviate from modeling the majority perspective.
+Appendix B.2 provides an overview of all metrics.
+First, we notice that the trends for and are both increasing—the first is expected, but the second requires an explanation.
+As the model is exposed to more annotations over the training iterations, the predicted label distribution starts to fit the true label distribution.
+However, here we consider each annotator individually: reports the average of the 10% lowest scores per annotator.
+The presence of disagreement implies the existence of annotators that annotate differently from the majority.
+Since our models predict the full distribution, they assign a proportional probability to dissenting annotators.
+Thus, learning to model the full distribution of annotations leads to an increase in .
+Second, we notice a difference between ACAL and AL.
+On MFTC and MHS, ACAL, compared to AL, yields overall smaller at the cost of a slower convergence in , showing the trade-off between modeling all annotators and representing minorities.
+However, with DICES the trend is the opposite.
+This is due to AL having access to the complete label distribution: it can model a balanced distribution, leading to lower worst-off performance.
+With a large number of annotations, ACAL requires more iterations to get the same balanced predicted distribution.
+Third, we observe differences among the annotator selection strategies ().
+shows the most differences—both and increase slower than for the other strategies.
+This suggests that selecting annotators based on the average embedding of the annotated content strongest emphasizes diverging label behavior.
+Finally, we analyze the impact of the sample selection strategies (, dotted vs.
+solid lines in Figure 3).
+For DICES, and lead to comparable results, likely due to the low number of samples.
+Using in MFTC leads to performance decreasing at the start of training.
+The strategy prioritizes obtaining annotations for already added samples to lower their entropy, while the variation in labels is irreconcilable (since there are limited labels available, and they are in disagreement).
+We see a similar pattern for MHS.
+These results further underline our main finding that ACAL is effective in representing diverse annotation perspectives when there is a (1) heterogeneous pool of annotators, and (2) a task that facilitates human label variation.
+5.4 Impact of subjectivity
+We further investigate ACAL strategies on (1) label entropy, and (2) cross-task performance.
+Alignment of ACAL strategies during training
+We want to investigate how well the ACAL strategies align with the overall subjective annotations: do they drive the model entropy in the right direction?
+We measure the entropy of the samples in the labeled training set at each iteration and compare it to the actual entropy of those samples.
+Higher entropy suggests that the selection strategy overestimates uncertainty.
+Lower entropy indicates that the model may not sufficiently account for disagreement.
+When the entropy matches the true entropy, the selection strategy is well-calibrated.
+We focus on DICES as a case study due to the wide range of entropy scores.
+We group each sample based on the true label entropy into low, medium, and high111Entropy bins: low (), medium () high ()..
+We apply the same categorization at each training iteration for samples labeled thus far.
+Subsequently, we plot the proportion of data points for which the selection strategy results in excessively high or excessively low entropy.
+[width=]tikz/align_training
+Figure 4 visualizes the proportions.
+At the beginning of training, entropy is generally low because samples have few annotations.
+Over time, the selected annotations better align with the true entropy.
+However, when and how much strategies succeeded in representing the true label distribution differs: and take longer to increase label entropy than the other two strategies.
+They are conservative in adding diverse labels.
+and increase the proportion of well-aligned data points earlier in the training process, achieving a balanced entropy alignment sooner.
+However, both strategies start to overshoot the target entropy, whereas the others show a more gradual alignment with the true entropy.
+This effect is strongest for .
+This finding suggests that minority-aware annotator-selection strategies achieve the best results in the early stages of training.
+They are effective for quickly raising entropy but can lead to overrepresentation.
+Cross-task performance
+Figure 5 compares the two annotator-centric metrics on the three tasks of MFTC and MHS—the datasets for which we have seen the least impact of ACAL over AL and PL.
+We select a data sampling () and annotator sampling strategy (), based on its strong performance on DICES for comprehensive comparison.
+[width=]tikz/comparison_tasks_SRTS
+When evaluating MFTC loyalty, which has the highest disagreement, is more accurately approximated with PL.
+Similarly, ACAL is outperformed by AL on for the dehumanize (high disagreement) task.
+However, for the less subjective task genocide, ACAL leads to higher .
+This suggests that the effectiveness of annotation strategies varies depending on the task’s degree of subjectivity and the available pool of annotators.
+The more heterogeneous the annotation behavior, indicative of a highly subjective task, the larger the pool of annotators required for each sample selection.
+We also observe that there is a trade-off between modeling the majority of annotators equally () and prioritizing the minority ().
 6 Conclusion
-We introduce Annotator-Centric Active Learning (ACAL), an active learning approach that incorporates annotator selection strategies aimed at capturing label variation among annotators.
-We experiment with tasks across three different datasets, each leading to different ACAL behaviors.
-One of these datasets, DICES, is the most realistic application of ACAL since the pool of possible annotators is the largest.
-Here, ACAL leads to more diverse label distributions using fewer annotations.
-However, we find that the effectiveness of the ACAL paradigm is contingent on data characteristics.
-These characteristics include the number of annotations per sample, the number of annotations per annotator, and the nature of disagreement in the task annotations.
-Our analysis shows that we can use these conditions to help explain the often disappointing results for AL in NLP applications.
-Including annotator-centric evaluation reveals how methods with similar averaged performance deal with different levels of disagreement among annotators.
-We show that evaluation can be enhanced by focusing on individual annotators, as there is a large gap between conventional, averaged, and worst-off performance.
-Furthermore, many aspects of our ACAL approach can be experimented with, e.g.
-by swapping the order in which samples are selected (in our case first) and annotators (second), or investigating the impact of including annotator-specific demographic information, as it is inconsistently predictive of annotation behavior Orlikowski et al.
-(2023); Beck et al.
+We present ACAL as an extension of AL to emphasize the selection of diverse annotators.
+We introduce three novel annotator selection strategies and four annotator-centric metrics and experiment with tasks across three different datasets.
+We find that the ACAL approach is especially effective in reducing the annotation budget when the pool of available annotators is large.
+However, its effectiveness is contingent on data characteristics such as the number of annotations per sample, the number of annotations per annotator, and the nature of disagreement in the task annotations.
+Furthermore, our novel evaluation metrics display the trade-off between modeling overall distributions of annotations and adequately accounting for minority voices, showing that different strategies can be tailored to meet different goals.
+Especially early in the training process, strategies that are aggressive in obtaining diverse labels have a beneficial impact Furthermore, we recognize that gathering a distribution that uniformly contains all possible (minority and majority) labels can be overly sensitive to small minorities or noise.
+Future work can integrate methods that account for noisy annotations Weber-Genzel et al.
+(2024) or that strike a balance between egalitarian and utilitarian approaches Lera-Leri et al.
 (2024).
 Limitations
-The main limitation of this work is that the experiments are based on simulated active learning which is known to bear potential issues (Margatina and Aletras, 2023).
+The main limitation of this work is that the experiments are based on simulated AL which is known to bear potential issues (Margatina and Aletras, 2023).
 In our study, a primary challenge arises with two of the datasets (MFTC, MHS), which, despite having a large pool of annotators, lack annotations from every annotator for each item.
@@
 Since we evaluate four different annotator selection strategies and two sample selection strategies across three datasets and seven tasks, the amount of experiments is high.
-This did not allow for further investigation of the difference using different classification models, the extensive turning of hyperparameters, or even different training paradigms.
+This did not allow for further investigation of other methods for measuring uncertainty (such as ensemble methods ), different classification models, the extensive turning of hyperparameters, or even different training paradigms (such as low-rank adaptation ).
 Lastly, a limitation of our annotator selection strategies is that they rely on a small annotation history.
@@
 Appendix A Detailed Experimental Setup
-A.1 Cross validation details
-We split the data on samples, meaning that all annotations for any given sample are completely contained in each separate split.
+Dataset | Task (dimension) | # Samples | # Annotators | # Annotations | # Annotations per item |
+DICES | Safety Judgment | 990 | 172 | 72,103 | 72.83 |
+MFTC | Morality (care) | 8,434 | 23 | 31,310 | 3.71 |
+MFTC | Morality (loyalty) | 3,288 | 23 | 12,803 | 3.89 |
+MFTC | Morality (betrayal) | 12,546 | 23 | 47,002 | 3.75 |
+MHS | Hate Speech (dehumanize, genocide, respect) | 17,282 | 7,807 | 57,980 | 3.35 |
+A.1 Dataset details
+We provide an overview of the datasets used in our work in Table A1. We split the data on samples, meaning that all annotations for any given sample are completely contained in each separate split.
 A.2 Hyperparameters
-We report the hyperparameters for training passive, AL, and ACAL in Tables 5, 6, and 7, respectively.
+We report the hyperparameters for training passive, AL, and ACAL in Tables A2, A3, and A4, respectively.
 For turning the learning rate for passive learning, on each dataset, we started with a learning rate of 1e-06 and increased it by a factor of 3 in steps until the model showed a tendency to overfit quickly (within a single epoch).
@@
 epochs per round | all | 20 |
-num rounds | all | 10 |
+num iterations | all | 10 |
 sample size | DICES | 79 |
@@
 learning rate | all | 1e-05 |
-num rounds | DICES | 50 |
-num rounds | MFTC (all), MHS (all) | 20 |
+num iterations | DICES | 50 |
+num iterations | MFTC (all), MHS (all) | 20 |
 epochs per round | DICES, MHS (all) | 20 |
@@
 uses a sentence embedding model to represent the content that an annotator has annotated.
-We use all-MiniLM-L6-v2111https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2.
+We use all-MiniLM-L6-v2222https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2.
 We select annotators that have not annotated yet (empty history) before picking from those with a history to prioritize filling the annotation history for each annotator.
@@
 A.5 Disagreement rates
-We report the average disagreement rates per dataset and task in Figure 5, for each of the dataset and task combinations.
+We report the average disagreement rates per dataset and task in Figure A1, for each of the dataset and task combinations.
+[width=]tikz/entropies
 Appendix B Detailed results overview
-B.1 Test set evaluation other MFTC and MHS tasks
-See Table 6 for the trade-off between data efficiency and test-set performance for the two conventional metrics ( and ).
-We include copy the earlier mentioned results for MFTC (care) and MHS (dehumanize) for convenience.
-B.2 Annotator-Centric evaluation for other MFTC and MHS tasks
-We show the full annotator-centric metrics results for MFTC betrayal (Table 8), MFTC loyalty (Table 9), MHS genocide (Table 10), and MHS respect (Table 11).
-Average | Worst-off | ||||
-App. | |||||
-0.578 | 0.147 | 0.420 | 0.199 | -1.6 | |
-0.581 | 0.149 | 0.433 | 0.212 | -1.6 | |
-0.593 | 0.161 | 0.430 | 0.239 | -5.0 | |
-0.583 | 0.148 | 0.429 | 0.199 | -1.6 | |
-0.594 | 0.150 | 0.419 | 0.203 | -2.5 | |
-0.584 | 0.148 | 0.434 | 0.200 | -1.3 | |
-0.588 | 0.149 | 0.435 | 0.204 | -1.0 | |
-0.591 | 0.149 | 0.428 | 0.194 | -2.6 | |
-0.589 | 0.147 | 0.431 | 0.195 | -48.6 | |
-0.589 | 0.149 | 0.430 | 0.200 | -0.0 | |
-passive | 0.481 | 0.199 | 0.360 | 0.290 | 0.0 |
-Average | Worst-off | ||||
-App. | |||||
-0.564 | 0.177 | 0.222 | 0.372 | -0.4 | |
-0.563 | 0.176 | 0.222 | 0.374 | -0.3 | |
-0.573 | 0.176 | 0.222 | 0.370 | -0.3 | |
-0.551 | 0.175 | 0.222 | 0.373 | -0.3 | |
-0.557 | 0.177 | 0.217 | 0.357 | -1.1 | |
-0.541 | 0.177 | 0.222 | 0.355 | -0.8 | |
-0.556 | 0.177 | 0.222 | 0.358 | -0.9 | |
-0.558 | 0.177 | 0.222 | 0.358 | -1.3 | |
-0.560 | 0.176 | 0.222 | 0.361 | -29.1 | |
-0.559 | 0.177 | 0.222 | 0.366 | -0.1 | |
-passive | 0.512 | 0.183 | 0.261 | 0.309 | 0.0 |
-Average | Worst-off | ||||
-App. | |||||
-0.700 | 0.227 | 0.000 | 0.560 | -6.3 | |
-0.698 | 0.225 | 0.000 | 0.565 | -1.7 | |
-0.700 | 0.224 | 0.000 | 0.566 | -1.7 | |
-0.702 | 0.224 | 0.000 | 0.565 | -1.7 | |
-0.711 | 0.229 | 0.000 | 0.549 | -12.6 | |
-0.707 | 0.231 | 0.000 | 0.548 | -7.9 | |
-0.709 | 0.231 | 0.000 | 0.548 | -7.9 | |
-0.712 | 0.229 | 0.000 | 0.547 | -12.6 | |
-0.339 | 0.387 | 0.000 | 0.496 | -60.1 | |
-0.331 | 0.390 | 0.000 | 0.497 | -24.7 | |
-passive | 0.700 | 0.245 | 0.000 | 0.570 | – |
-Average | Worst-off | ||||
-App. | |||||
-0.460 | 0.331 | 0.000 | 0.528 | -18.8 | |
-0.456 | 0.331 | 0.000 | 0.530 | -18.8 | |
-0.461 | 0.331 | 0.000 | 0.529 | -18.8 | |
-0.460 | 0.331 | 0.000 | 0.528 | -18.8 | |
-0.466 | 0.323 | 0.000 | 0.533 | -4.9 | |
-0.463 | 0.323 | 0.000 | 0.532 | -4.9 | |
-0.459 | 0.324 | 0.000 | 0.531 | -4.9 | |
-0.462 | 0.324 | 0.000 | 0.532 | -4.9 | |
-0.339 | 0.387 | 0.000 | 0.496 | -60.1 | |
-0.331 | 0.390 | 0.000 | 0.497 | -24.7 | |
-passive | 0.259 | 0.405 | 0.000 | 0.587 | – |
-B.3 Training process
+B.1 Annotator-Centric evaluation for other MFTC and MHS tasks
+We show the full annotator-centric metrics results for MFTC betrayal, MFTC loyalty, MHS genocide, and MHS respect in Table B1.
+This follows the same format at Table 1.
+The results in this table also form the basis for Figure 5.
+Average | Worst-off | |||||||
+App. | ||||||||
+MFTC (betrayal) | 71.5 | .047 | 57.8 | .147 | 42.0 | .199 | -1.6 | |
+71.2 | .046 | 58.1 | .149 | 43.3 | .212 | -1.6 | ||
+71.2 | .051 | 59.3 | .161 | 43.0 | .239 | -5.0 | ||
+71.0 | .046 | 58.3 | .148 | 42.9 | .199 | -1.6 | ||
+72.6 | .042 | 59.4 | .150 | 41.9 | .203 | -2.5 | ||
+73.6 | .045 | 58.4 | .148 | 43.4 | .200 | -1.3 | ||
+74.0 | .045 | 58.8 | .149 | 43.5 | .204 | -1.0 | ||
+73.2 | .044 | 59.1 | .149 | 42.8 | .194 | -2.6 | ||
+72.1 | .046 | 58.9 | .147 | 43.1 | .195 | -48.6 | ||
+71.8 | .047 | 58.9 | .149 | 43.0 | .200 | -0.0 | ||
+PL | 75.2 | .037 | 48.1 | .199 | 36.0 | .290 | 0.0 | |
+MFTC (betrayal) | 66.9 | .034 | 56.4 | .177 | 22.2 | .372 | -0.4 | |
+68.9 | .032 | 56.3 | .176 | 22.2 | .374 | -0.3 | ||
+67.1 | .031 | 57.3 | .176 | 22.2 | .370 | -0.3 | ||
+68.4 | .031 | 55.1 | .175 | 22.2 | .373 | -0.3 | ||
+61.3 | .032 | 55.7 | .177 | 21.7 | .357 | -1.1 | ||
+66.5 | .032 | 54.1 | .177 | 22.2 | .355 | -0.8 | ||
+62.4 | .033 | 55.6 | .177 | 22.2 | .358 | -0.9 | ||
+64.4 | .031 | 55.8 | .177 | 22.2 | .358 | -1.3 | ||
+71.5 | .030 | 56.0 | .176 | 22.2 | .361 | -29.1 | ||
+66.5 | .033 | 55.9 | .177 | 22.2 | .366 | -0.1 | ||
+PL | 62.5 | .029 | 51.2 | .183 | 26.1 | .309 | 0.0 | |
+MHS (genocide) | 26.5 | .050 | 70.0 | .227 | 0.0 | .560 | -6.3 | |
+28.2 | .051 | 69.8 | .225 | 0.0 | .565 | -1.7 | ||
+28.1 | .051 | 70.0 | .224 | 0.0 | .566 | -1.7 | ||
+28.3 | .050 | 70.2 | .224 | 0.0 | .565 | -1.7 | ||
+32.8 | .077 | 71.1 | .229 | 0.0 | .549 | -12.6 | ||
+27.7 | .048 | 70.7 | .231 | 0.0 | .548 | -7.9 | ||
+26.7 | .048 | 70.9 | .231 | 0.0 | .548 | -7.9 | ||
+27.3 | .048 | 71.2 | .229 | 0.0 | .547 | -12.6 | ||
+28.0 | .048 | 33.9 | .387 | 0.0 | .496 | -60.1 | ||
+33.3 | .080 | 33.1 | .390 | 0.0 | .497 | -24.7 | ||
+PL | 21.6 | .044 | 70.0 | .245 | 0.0 | .570 | – | |
+MHS (respect) | 41.4 | .086 | 46.0 | .331 | 0.0 | .528 | -18.8 | |
+40.8 | .087 | 45.6 | .331 | 0.0 | .530 | -18.8 | ||
+41.2 | .086 | 46.1 | .331 | 0.0 | .529 | -18.8 | ||
+40.6 | .086 | 46.0 | .331 | 0.0 | .528 | -18.8 | ||
+32.8 | .077 | 46.6 | .323 | 0.0 | .533 | -4.9 | ||
+41.0 | .085 | 46.3 | .323 | 0.0 | .532 | -4.9 | ||
+41.8 | .084 | 45.9 | .324 | 0.0 | .531 | -4.9 | ||
+40.6 | .085 | 46.2 | .324 | 0.0 | .532 | -4.9 | ||
+41.7 | .085 | 33.9 | .387 | 0.0 | .496 | -60.1 | ||
+33.3 | .080 | 33.1 | .390 | 0.0 | .497 | -24.7 | ||
+PL | 41.0 | .080 | 25.9 | .405 | 0.0 | .587 | – |
+B.2 Training process
 In our main paper, we report a condensed version of all metrics during the training phase of the active learning approaches.
 Below, we provide a complete overview of all approaches over all metrics.
-The results can be seen in Figures 7 through 13.
+The results can be seen in Figures B1 through B7.
+[width=]tikz/training_process_DICES_overall
+[width=]tikz/training_process_MFTC_care
+[width=]tikz/training_process_MFTC_loyalty
+[width=]tikz/training_process_MFTC_betrayal
+[width=]tikz/training_process_MHS_dehumanize
+[width=]tikz/training_process_MHS_genocide
+[width=]tikz/training_process_MHS_respect